When SrF2, strontium fluoride, is added to water, the salt dissolves to a very small extent according to the reaction below. At equilibrium the concentration of Sr2+ is found to be 0.00196 M. What is the value of Ksp for SrF2? (Please give your answer with 2 significant figures.)




1
Expert's answer
2021-09-16T02:01:09-0400

"[Sr^{2+}] = 0.00196 \;M = S"

SrF2 <=> Sr2+(aq) + 2F-(aq)

"k_{sp} = [Sr^{+2}][F^-]^2 \\

k_{sp} = (0.00196)(2 \times 0.00196)^2 \\

k_{sp} = 3.0 \times 10^{-8}"